### v1

# Diffusion-Refined VQA Annotations for Semi-Supervised Gaze Following

Abstract
Training gaze following models requires a large number of images with gaze target coordinates annotated by human annotators, which is a laborious and inherently ambiguous process. We propose the first semi-supervised method for gaze following by introducing two novel priors to the task. We obtain the first prior using a large pretrained Visual Question Answering (VQA) model, where we compute Grad-CAM heatmaps by ‘prompting’ the VQA model with a gaze following question. These heatmaps can be noisy and not suited for use in training. The need to refine these noisy annotations leads us to incorporate a second prior. We utilize a diffusion model trained on limited human annotations and modify the reverse sampling process to refine the Grad-CAM heatmaps. By tuning the diffusion process we achieve a trade-off between the human annotation prior and the VQA heatmap prior, which retains the useful VQA prior information while exhibiting similar properties to the training data distribution. Our method outperforms simple pseudo-annotation generation baselines on the GazeFollow image dataset. More importantly, our pseudo-annotation strategy, applied to a widely used supervised gaze following model (VAT), reduces the annotation need by 50%. Our method also performs the best on the VideoAttentionTarget dataset.
Keywords:
Gaze following Semi-supervised learning Diffusion Model
1 Introduction
Human gaze behavior is a vital cue for applications such as human-machine interaction [39, 11, 1], social interaction analysis [35, 5], and human intention interpretation [48, 52]. Although eye-tracking glasses can be used to track gaze behavior, they are intrusive, require personalized calibration, and are mostly restricted to laboratory environments. In contrast, the gaze following task [45, 9, 12] predicts a person’s gaze target in an image in-the-wild, by predicting a target heatmap from an input scene image with the person’s head bounding box.
The effectiveness of gaze following methods depends on the quantity and quality of training data with annotated gaze targets. Annotating gaze targets is a laborious and ambiguous process. For an image and a subject in the image, an annotator must consider the subject’s head orientation, gaze angle, and follow the gaze direction to determine where the point of gaze falls on an object or surface within the image. This can be very challenging, as the subject may be looking away from the camera, their eyes may be occluded, the image may be cluttered, and there may be multiple plausible targets, as shown in Fig. 1(a).
Figure 1: (a) Gaze following annotation challenges. Annotating gaze is a laborious task with inherent ambiguities. (b) Pseudo annotations for gaze following. We generate pseudo annotations by first computing Grad-CAM heatmaps from a pre-trained VQA model, and then refining the noisy heatmaps with a diffusion model.
It is much more efficient to utilize large quantities of unlabeled data, along with a limited amount of labeled data, and leverage semi-supervised learning to improve model performance. Semi-supervised learning has been successful in tasks such as image recognition and semantic segmentation [64, 59, 26, 2, 23]. In the context of gaze following, semi-supervised learning can be particularly useful when developing models for a specific scene, such as a psychology lab conducting social interaction studies or a supermarket studying customer gaze behaviors, by requiring only a small fraction of the collected data to be annotated.
This is the first semi-supervised gaze following method, to the best of our knowledge. Semi-supervised recognition and segmentation methods are typically not directly applicable to gaze following due to task differences. As gaze following models predict a target heatmap instead of classification logits and require an intact scene image as input, methods that operate on the predicted multi-class distribution [57, 26], learn disentangled representations [42, 33], or perform strong data augmentation (e.g.,CutOut[54, 13]) cannot easily be adapted to gaze following. Additionally, for semi-supervised methods that use a teacher model to pseudo-annotate the unlabeled data [30, 59, 29], it is unclear how to design a gaze following teacher model to generate good quality pseudo-annotations.
To generate high-quality pseudo-gaze labels, our semi-supervised method combines the power of pre-trained large vision-language (VL) models with an annotation refinement method. Due to their extensive training sets, VL models have naturally acquired a wealth of knowledge, including the knowledge for inferring human gaze targets. Specifically, we “prompt” a pre-trained Visual Question Answering (VQA) model [37, 44, 67] with an appropriate question to obtain the text description of the gaze target of a subject in the image. We then use Grad-CAM [49, 6], to ground the VQA model’s description to spatial locations, referred to as Grad-CAM heatmaps. We want to turn these heatmaps into pseudo-annotations for training gaze following models.
There are technical challenges in utilizing Grad-CAM heatmaps. First, Grad-CAM heatmaps do not precisely overlap with the gaze targets because they are dispersed and noisy, as shown in Fig. 1(b). Second, some Grad-CAM heatmaps do not provide useful prior knowledge and may highlight incorrect gaze target locations. Hence, the noisy Grad-CAM heatmaps should be refined into clean pseudo-annotations, with a trade-off between retaining the Grad-CAM heatmap information and the pre-trained prior on the annotations available.
We choose a diffusion model as the annotation prior of our refinement method. We are inspired by the recent use of diffusion models as inverse problem solvers and image editors that modify the reverse sampling process with a conditional input [36, 56, 41]. The goal is to generate an output image that retains the semantics of the degraded input (here the Grad-CAM heatmap), while having similar properties (i.e., geometry and gaze context) to the training data distribution.
We begin by training a diffusion model to capture the distribution of human-labeled annotations. The trained diffusion model then generates refined annotations by running the reverse sampling process, initialized from Grad-CAM heatmaps with an appropriate level of Gaussian noise. The added noise serves to smooth artifacts and noisy activations in the Grad-CAM heatmaps; the magnitude of the noise dictates how much of the information is preserved.
In summary, we propose the first semi-supervised gaze following method, in which we introduce two novel priors to the task: (1) Prior knowledge from pre-trained vision-language models for initial annotation generation. (2) A novel diffusion-based annotation prior to refine the noisy VL annotations into reliable pseudo-labels.
Our method outperforms pseudo-annotation generation baselines on the GazeFollow image dataset [45]. More importantly, our pseudo-annotation strategy, applied to the widely used VAT model [9], outperforms the fully supervised model trained with double the amount of annotations when only 5% and 10% labels are available. Our method also performs the best on the VideoAttentionTarget dataset [9], where we adapt a pre-trained gaze following model to new videos using only around 100 (2–23%) annotated frames.
2 Prior Work
Gaze Following was first introduced in [45] together with the GazeFollow dataset and a model composed of two separate pathways for encoding gaze orientation and scene saliency information, also adopted by later work [34, 8, 9, 12]. Chong et al. [8] considered out-of-frame targets and also extended the task to videos [9], proposing the VideoAttentionTarget dataset and the VAT model. Later work has improved gaze following models by leveraging monocular depth estimations [12, 25, 60, 21] and human poses [16, 4]. Tu et al. [62] proposed a transformer model for gaze following, while [24, 38] added training losses for numerical coordinate regression and patch-level gaze distribution prediction. Recent works investigate 3D gaze following [22, 20], gaze following for children [58], and object-aware gaze target prediction [61]. All of these methods are fully-supervised.
Semi-supervised Learning. The majority of semi-supervised learning methods adopt a teacher-student framework, which can be categorized into two groups: self-training [30, 7, 15, 54] and consistency regularization [29, 59, 13, 18]. Self-training methods generate pseudo annotations from teacher models trained with labeled data, while consistency regularization applies a consistency loss between the student and teacher model outputs, updating the teacher model gradually during training. As most semi-supervised learning methods are built for recognition or segmentation tasks, they are not directly applicable to gaze following [54, 42, 13, 26] due to task differences. However, some general methods that do not perform task-specific operations [29, 59] can be applied to gaze following with appropriate modifications (details in the Supplementary). In our method, we adopt a self-training pipeline using the VQA prior and diffusion-based annotation prior, and show that we can get further improvements by using applicable consistency regularization methods (e.g. Mean Teacher [59]) to enhance the training of the diffusion model.
Vision-Language Models. Vision-Language (VL) models [63, 27, 51, 31] have succeeded in tasks such as visual grounding, visual question answering (VQA), and image captioning. They also do well in low-shot generalization tasks by “prompting” [43, 37, 31], where appropriate task instructions (usually in the form of text) are given to a pretrained model to generate outputs useful for other tasks. Grad-CAM [50] visualizations on the cross-attention maps localize the linguistic input effect on the visual input [32, 66] thus achieving a degree of interpretability.
Diffusion Models. Denoising diffusion probabilistic models (DDPM) [53, 19] give state-of-the-art results on image synthesis, by capturing intricate data properties [10, 40, 46, 28]. They have also been used in inverse problems [56], image editing [36], and adversarial purification [41]. These works treat images with added noise as intermediate steps of the reverse sampling process, showing that the noise magnitude affects the information retained from input images. In our task, we add appropriate magnitude noise to the Grad-CAM heatmaps to achieve the best trade-off between the VQA prior and annotation prior. Recent work adopts diffusion models in semi-supervised learning. [18] uses a diffusion model as the student model for 3D object detection, whereas [65] trains a diffusion model for image generation with images pseudo-labeled by existing semi-supervised models. In contrast, we use diffusion model as annotation prior, to refine the initial noisy annotations from VQA into high-quality pseudo annotations.
3 Proposed Method
3.1 Overview
Our approach is illustrated in Fig. 2a. The goal is to leverage the knowledge priors from a VQA model to improve gaze following with minimal manual annotation. To achieve this, we utilize Grad-CAM heatmaps of a VQA model that is prompted with gaze-related questions.
However, the Grad-CAM heatmaps need refinement to be useful for training supervision. For this, we train a diffusion model on the groundtruth heatmaps from labeled data to learn to generate heatmaps from the human-label distribution. We use the model to refine the “noisy” Grad-CAM into high-quality pseudo annotations. We train existing gaze following models on both groundtruth heatmaps from labeled data and refined heatmaps from unlabeled data.
Suppose we only have a small number of images that are labeled with gaze target locations =${\mathcal{L}=\{(\bm{X}_{i},\bm{p}_{i})\}_{i=1}^{N_{l}}}$ and a larger number of unlabeled images ${\mathcal{U}=\{\bm{X}_{j}\}_{j=1}^{N_{u}}}$. $\bm{X}_{i}$ indicates an input triplet to the gaze following models, which includes a scene image $\bm{I}_{i}\in R^{3\times H\times W}$, the cropped head image $\bm{I}_{h}^{i}\in R^{3\times H\times W}$, and the head location binary mask $\bm{M}_{h}^{i}\in R^{H\times W}$. We represent the annotated gaze coordinate of the "gazing" person as $\bm{p}_{i}=[p_{i}^{x},p_{i}^{y}]$. For each $\bm{p}_{i}$, a ground truth heatmap $\mathbf{h}^{i}\in R^{H^{\prime}\times W^{\prime}}$ is generated by placing a Gaussian of fixed variance centered at ${\bm{p}_{i}}$, as is done in previous gaze following frameworks [9, 12, 34].
Figure 2: (a) Overall pipeline for semi-supervised training. We compute Grad-CAM heatmaps for a large set of unlabeled images and train the diffusion model with a small human-labeled set. We use the diffusion model to refine the Grad-CAM heatmaps into pseudo-annotations. Both the pseudo-annotations and the human-labeled training set are used to train a gaze following model. (b) Grad-CAM heatmap generation procedure. Given an input image with an overlaid bounding box for the person, we “prompt" a large pretrained VQA model with a gaze question and compute the Grad-CAM heatmap from the answer. (c) Grad-CAM refinement process. We perturb the raw Grad-CAM heatmaps with Gaussian noise and pass through the reverse diffusion process to generate the refined pseudo-annotations.
3.2 Grad-CAM heatmap extraction
We use OFA [63], an open-source large pretrained VL model as the VQA model. Fig. 2b illustrates the procedure for Grad-CAM heatmap generation. Given an image $\bm{I}$ and the head bounding box of a person $\bm{l}_{h}$, we use Mask R-CNN [17] for person detection and find the bounding box that maximally overlaps with $\bm{l}_{h}$. We overlay this bounding box on top of the input image and give it as input to the VQA model along with a “prompt” in the form of a question: “What is the person in the green bounding box looking at?".
OFA has a transformer-based encoder-decoder structure. The model generates a sequence of words as an answer. We compute a Grad-CAM heatmap $\mathcal{G}=\{\bm{g}_{j}\}_{j=1}^{N_{u}}$ on the decoder cross-attention weights between the input query token, and the image patch tokens based on [6] after selecting the noun from the answer. Details are in the supplementary.
In our experiments, we found that Grad-CAM heatmaps are beneficial for gaze following. Naively using these heatmaps as additional input to the current gaze following model results in a significant performance boost. However, in real applications, computing Grad-CAM heatmaps at test-time can be impractical due to the memory and time costs of accessing large VL models. Instead we use Grad-CAM heatmaps offline as pseudo-annotations for training a gaze following model. As discussed, using noisy Grad-CAMs directly as pseudo labels can hurt final model performance. Thus, we propose to utilize a diffusion model to refine the initial Grad-CAM heatmaps into more suitable pseudo labels for training.
3.3 Diffusion Model Training
As there is no available diffusion model for the gaze following task, we first train a diffusion model on the labeled data to generate gaze heatmaps $\mathbf{h}\sim p(\mathbf{h|\bm{X}})$, following the general training procedure of DDPM [19]. For ease of understanding, we provide a synopsis of Gaussian diffusion models.
Figure 3: Diffusion model training and refinement. (a) The diffusion model is trained on supervised data with noise added to the ground truth heatmap at random time steps. (b) During refinement, we add noise at a specific time step to the Grad-CAM heatmap. We treat this heatmap as an intermediate step input during the reverse process. Heatmaps are overlayed on the original images for illustration purposes. The conditional feature extraction for the diffusion model is omitted for simplicity.
The diffusion model consists of a forward process that gradually corrupts the input by adding Gaussian noise, and a reverse process that iteratively denoises the noisy input. The forward process is defined by a noise schedule $\alpha_{t}$ as:
$q(\mathbf{h}_{t}|\mathbf{h}_{t-1})$ $:=\mathcal{N}(\mathbf{h}_{t};\sqrt{\alpha_{t}}\mathbf{h}_{t-1},(1-\alpha_{t})\mathbf{I})$ (1)
$q(\mathbf{h}_{1:T}|\mathbf{h}_{0})$ $:=\prod_{t=1}^{T}q(\mathbf{h}_{t-1}|\mathbf{h}_{t}).$ (2)
where $\mathbf{h}_{0}$ is the ground truth heatmap $\mathbf{h}$, $\mathbf{h}_{1},\dots,\mathbf{h}_{T-1}$ are intermediate latent variables that represent noisy versions of $\mathbf{h}_{0}$, and $\mathbf{h}_{T}$ is the terminal state which corresponds to a unit Gaussian distribution: $\mathbf{h}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
The reverse process uses Gaussian transitions with fixed covariance $\sigma_{t}$ and gradually denoises the data, starting from $\mathbf{h}_{T}$. Since $q(\mathbf{h}_{t-1}|\mathbf{h}_{t})$ is intractable, it is approximated with a neural network:
$p_{\theta}(\mathbf{h}_{t-1}\mid\mathbf{h}_{t}):=\mathcal{N}(\mathbf{h}_{t-1};\bm{\mu}_{\theta}(\mathbf{h}_{t},t),\sigma_{t}\mathbf{I}).$ (3)
A useful property of the forward process is that it allows sampling of any intermediate $\mathbf{h}_{t}$ given $\mathbf{h}_{0}$, and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$:
$q(\mathbf{h}_{t}|\mathbf{h}_{0})=\mathcal{N}(\mathbf{h}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{h}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$ (4)
which is utilized in learning the reverse process. For that, we adopt the $\mathbf{x}_{0}$-parameterization as in [3, 28] and directly predict the final $\mathbf{h}_{0}$, leading to the following minimization objective, where $\mathbf{h}_{0}^{\theta}$ is parameterized by a U-Net [47] :
$\mathbb{E}_{\mathbf{h}_{0},t}\left[\lVert\mathbf{h}_{0}-\mathbf{h}_{0}^{\theta}(\mathbf{h}_{t},t)\rVert^{2}\right],$ (5)
Fig. 3(a) shows the training procedure. At each iteration, we sample a random time step $t$ from $U[1,T]$ to generate a noisy input $\mathbf{h}_{t}$ using Eq. 4. We feed $\mathbf{h}_{t}$ to the diffusion model for single-step denoising, and optimize with the loss of Eq. 5. We condition the diffusion model with the gaze feature $\mathbf{c}$ extracted from the input triplet $\bm{X}$ (input image, head crop, and head location mask) to output a gaze heatmap grounded on the gazing person; the parametrization of the U-Net denoiser becomes $\mathbf{h}_{0}^{\theta}(\mathbf{h}_{t},t,\mathbf{c})$. The feature extractor follows the structure of the VAT model [9], which takes the input triplet $\bm{X}$ and outputs the extracted features concatenated from two pathways for encoding the scene and gaze features. Details of the VAT structure are shown in Supplementary. Following [14], we add the extracted features as conditional features to each layer in the U-Net denoiser with the necessary transposed convolution layers to match the feature size of U-Net. During the reverse sampling process, we used DDIM [55] to sample the next step input using the predicted $\mathbf{h}_{0}$ from the U-Net.
3.4 Heatmap Refinement using Diffusion Models
After training on labeled images $\mathcal{L}$, the learned diffusion process models an annotation prior over ground truth annotations $p(\mathbf{h|\bm{X}})$. During the semi-supervised training process, the trained diffusion model is applied to the unsupervised data $\mathcal{U}$ to produce the refined pseudo annotations.
The refinement process is shown in Fig. 3(b). Given a noisy Grad-CAM heatmap $\bm{g}$ of an unsupervised input sample $\bm{X}$, we first add an appropriate amount of noise using the forward diffusion process $\hat{\mathbf{h}}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{g}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$ according to Eq. 4. The added noise smoothes out the noise artifacts of the Grad-CAM heatmaps, while preserving the information of the highlighted regions. Then, by iteratively applying the steps of the denoising procedure in Eq. 3, we get samples of the final heatmap $\hat{\mathbf{h}}_{0}$ that are geometrically and contextually similar to the human-labeled heatmap $\mathbf{h}_{0}$ while retaining gaze target information from the noisy Grad-CAM outputs.
The amount of information retained depends on the choice of $t$ (i.e., the magnitude of added noise), as proven in [36, 41]. A larger magnitude of noise corrupts the Grad-CAM heatmaps more, decreasing the similarity between the diffusion output and the Grad-CAM priors. Similar to [36] which achieves a good trade-off between realism and faithfulness in image editing with an intermediate level of noise, we also found that initializing from an intermediate timestep in the diffusion process achieved the best trade-off between the Grad-CAM information and the model’s learned distribution $p(\mathbf{h|\bm{X}})$. With this formulation, the diffusion model can incorporate the high-quality Grad-CAM heatmaps, while ignoring the Grad-CAM priors when they are highly noisy or highlight unlikely gaze target locations. We provide analyses of the added noise magnitude in Sec. 4.3.2.
3.5 Semi-supervised Training
In semi-supervised training, we use the refined heatmaps $\mathcal{R}$ as pseudo labels for the unsupervised data $\mathcal{U}$. We train the gaze following model with the set of ground-truth heatmaps $\mathcal{H}$ from the labeled data $\mathcal{L}$ and the refined heatmaps $\mathcal{R}$ from the unlabeled data $\mathcal{U}$. We use the Mean Squared Error (MSE) loss between the predicted heatmap $\bm{h^{\prime}}$ and the target label ${\bf h}$. ${\bf h}$ is either a human annotation (labeled data) or a pseudo annotation (unlabeled data). We minimize:
$\mathcal{L}=\frac{1}{|\mathcal{H}\cup\mathcal{R}|}\sum_{{\bf h}\in\mathcal{H}\cup\mathcal{R}}\mathcal{L}_{mse}(\bm{h^{\prime}},\bm{h}).$ (6)
4 Experiments
4.1 Setting and Implementation Details
Datasets. GazeFollow [45], the largest real image dataset, contains the annotated gaze targets of 130,339 people in 122,143 images. 4782 people are used for testing, and the rest are used in training. For the test set, there are 10 annotations per person-image pair to account for the ambiguity of the gaze target, whereas the training set contains a single annotation for each pair. VideoAttentionTarget [9] is a video dataset of 50 different shows collected from YouTube. Each person-image pair in both training and test set has only one annotation.
Evaluation metrics. We employ evaluation metrics suggested by previous work [45, 9, 34]. The distance metric (Dist.) refers to the normalized $L_{2}$ distance between the predicted gaze point (point with the highest heatmap response) and the ground truth location. In GazeFollow reports both average (AvgDist) and minimum distances (MinDist). Area Under Curve (AUC) evaluates the concordance of predicted heatmaps with ground truth [9]. For GazeFollow, ground truth is the 10 annotations for each test set image, whereas in VideoAttentionTarget it is a thresholded Gaussian centered at the single given annotation.
Implementation Details. We used a diffusion model with a linear noise schedule and $T{=}500$ steps. The heatmap size was $64{\times}64$. We used a batch size of 80 and a learning rate of $2.5\times 10^{-4}$ in the semi-supervised learning experiments. For training the diffusion model on the pool of labeled data, we used a learning rate of $5\times 10^{-5}$. When refining Grad-CAM heatmaps, we used DDIM [55] for inference, with two inference steps.
4.2 Semi-supervised Training Results
We build our gaze following experiments around the popular, publicly available VAT model [9]. To showcase the impact of our pseudo annotation generation method, we did not use additional modalities, processing, or supervisions, such as depth input, pose estimations, and object-level annotations [12, 4, 16, 61] in all experiments for the baseline methods and the diffusion model.
Following the standard experimental settings in semi-supervised learning [13, 30, 59], we considered different amounts of annotations from the GazeFollow training set, namely 5%, 10%, and 20%, and treated the remaining data as unlabeled. The trained student model is evaluated on the test set in all cases.
We used VAT as the student model and compared our method of generating pseudo annotations with several alternatives [^1] Implementation details of the baselines are provided in the supplementary material.: 1) Semi-VAT: the pseudo-annotations for the unlabeled data were generated by the VAT model trained with labeled images. 2) Semi-VAT-GC: We created VAT-GC by adding Grad-CAM heatmaps as conditional inputs to the VAT model and modifying the VAT model accordingly. VAT-GC was trained on the labeled data to generate pseudo annotations for the unlabeled data. 3) VAT-MT: We use the Mean Teacher (MT) method [59] as a representative consistency regularization method to train the VAT model, which updates the teacher model using both labeled and unlabeled data through an exponential moving average (EMA) of student model weights. Here we applied the consistency loss in MT between the heatmap output from the teacher and student models. The first two baselines are self-training methods, while the 3rd belongs to the consistency regularization category.
On the other hand, we also build two versions of our method: 1) GradCAM-Diffusion-Refinement (GCDR): our method of using a diffusion model trained with labeled data to refine the ‘noisy’ Grad-CAM heatmaps and use them as pseudo-annotations. 2) GCDR-MT: we used the Mean-Teacher method to train the diffusion model with both labeled and unlabeled data, and this enhanced diffusion model was used to refine the Grad-CAM heatmaps.
Table 1: Results of semi-supervised training methods with different ratios of labeled data on GazeFollow dataset. In the top and bottom rows, we show the performances of VAT trained with supervised data only or with full training data, to show the potential lower and upper limits. Best numbers are marked as bold. Our methods outperform supervised VAT trained with double the amount of annotations.
  Method 5% labels 10% labels 20% labels
Dist. $\downarrow$ AUC $\uparrow$ Dist $\downarrow$ AUC $\uparrow$ Dist. $\downarrow$ AUC $\uparrow$
Avg. Min. Avg. Min. Avg. Min.
  VAT (Supervised) 0.230 0.161 0.835 0.202 0.133 0.869 0.182 0.116 0.892
  Semi-VAT 0.222 0.152 0.846 0.195 0.128 0.875 0.176 0.110 0.896
Semi-VAT-GC 0.217 0.149 0.846 0.195 0.127 0.879 0.178 0.112 0.895
VAT-MT 0.219 0.150 0.850 0.189 0.122 0.882 0.174 0.108 0.898
GCDR (Ours) 0.201 0.135 0.863 0.179 0.115 0.886 0.166 0.103 0.902
GCDR-MT (Ours) 0.194 0.128 0.870 0.172 0.108 0.892 0.162 0.098 0.904
  VAT(100% labels) [9] 0.137 0.077 0.921
Semi-supervised training results are in Tab. 1. Our two proposed methods consistently outperform the three semi-supervised baselines in all three annotation scenarios. When training with 5% and 10% annotations, the performance of GCDR-MT surpasses the supervised VAT trained with double the amount of annotations. Results of training with 50% are in the supplementary.
Our methods show more prominent improvements in Dist. compared to AUC. The AUC on GazeFollow evaluates the concordance of the predicted heatmap with the 10 annotations on each test image, while Dist. evaluates the $L_{2}$ distance between the predicted gaze point and the annotations. Therefore, the results mean that we did better in predicting the probable target location than predicting the exact shape of the heatmap. The refined heatmaps from the diffusion model closely resemble the spatial distribution of a human-labeled annotation, which is a single Gaussian (Fig. 4). In contrast, the baseline methods predict the target location less accurately, while outputting less certain and larger heatmaps which tend to overlap more with the group-level annotations, thus favoring the AUC (see supplementary material).
VAT VAT-GC Raw Grad-CAM GCDR Groundtruth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Visualizations of pseudo heatmaps generated by different teachers. Our method generates the cleanest pseudo annotations while retaining the Grad-CAM heatmaps priors (Rows 1–3). When the initial Grad-CAM heatmap responds strongly to unlikely locations or is completely noisy, our method can also ignore it (Row 4).
Fig. 4 visualizes the pseudo annotations predicted by the teacher models. In the top three rows, our model retains positional priors from the Grad-CAM heatmaps with almost identical structure to ground truth heatmaps. On the other hand, when the raw Grad-CAM heatmaps are inaccurate or very noisy (Row 4), our method can ignore the Grad-CAM heatmaps.
4.3 Ablation Studies
In the ablation studies shown in Tab. 3, we tested the contribution of the various components of our method. We also tested different parameters of the diffusion sampling process in Tab. 3. All ablation studies were performed using 10% of the GazeFollow labels and the teacher models were trained only on supervised data (without MT) to simplify training. Additionally, in Tab. 4, we analyze the effect of the VQA priors by training the baselines in fully-supervised settings with/without Grad-CAM heatmaps.
Table 2: Ablations of ways for generating the pseudo annotations.
  Method Dist. $\downarrow$ AUC $\uparrow$
Avg. Min.
  No Grad-CAM 0.191 0.125 0.867
No Refine 0.237 0.166 0.833
Argmax Refine 0.207 0.139 0.869
Direct Mapping 0.190 0.122 0.878
  Proposed GCDR 0.179 0.115 0.886
Table 3: Ablations of parameters of the diffusion model refining process.
  Noise.$t$ Inf. Steps Dist. $\downarrow$ AUC $\uparrow$
Avg. Min.
  300 2 0.186 0.120 0.877
200 2 0.184 0.117 0.883
250 5 0.180 0.115 0.883
250 2 0.179 0.115 0.886
4.3.1 Refinement methods
In No Grad-CAM, the diffusion model samples pseudo annotations from Gaussian noise without using Grad-CAM heatmaps. Without the VQA model priors, there is a significant performance decrease. No Refinement directly uses the raw Grad-CAM heatmaps as pseudo annotations without refinement, which also leads to a significant performance decrease due to the noisy nature of the Grad-CAM heatmaps. Argmax Refinement generates a Gaussian heatmap from the maximum point of the Grad-CAM heatmap as pseudo annotation. Despite the improvement over No refinement due to the “cleaner" annotation pattern, performance is still far behind our method, because some of the Grad-CAM heatmaps are low-quality (Fig. 4), so Gaussian ends up on outliers or incorrect locations. On the contrary, our method can ignore the Grad-CAM heatmaps in these cases. In Direct Mapping, we trained a U-Net to learn a “direct mapping” from Grad-CAM heatmaps to groundtruth heatmaps which we then applied to the unlabeled data. This shows a large performance decrease, because it cannot capture the full distribution that a diffusion model learns.
All the above ablations validate the importance of both the Grad-CAM heatmaps priors and the priors that the diffusion model learns.
4.3.2 Added Noise Magnitude Effects
In this section, we analyze the effect of the added noise magnitude on the diffusion output. As shown in Tab. 3, adding noise at the 250th step achieved the best trade-off between the VQA prior and the pretrained annotation prior. As a reminder, the later the timestep we add the noise, the larger its magnitude (the more the model ignores the Grad-CAM).
Grad-CAM $t=100$ $\bm{t=250}$ $t=400$ Pure noise
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Diffusion model output for noise added at different timesteps. Red dots represent the ground truth annotation. Adding noise at earlier steps generates outputs on high Grad-CAM response regions. Adding noise at later steps generates outputs similar to sampling from pure noise. Noise at step 250 is the best trade-off.
We further visualized the outputs from the same diffusion model when adding noise at different time steps in Fig. 5. In the top 2 rows, the Grad-CAM heatmaps highlight correct locations, whereas the Grad-CAM heatmaps of the bottom 2 rows are noisy or highlight incorrect locations. In all these cases, when adding noise at the 100th step, the outputs are located on the highlighted region of the Grad-CAM heatmap. In contrast, when adding noise at the 400th step, the large magnitude of added noise leads to outputs similar to sampling directly from Gaussian noise without using the Grad-CAM heatmaps. Adding noise at step 250 appears to be the best trade-off.
Table 4: Results of using Grad-CAM heatmaps as a direct input. When Grad-CAM heatmaps are directly input into the gaze inference process, both the VAT and the diffusion model show significant improvement in performance.
  Method Auxiliary Input 5% labels 10% labels 20% labels
Dist. $\downarrow$ AUC $\uparrow$ Dist $\downarrow$ AUC $\uparrow$ Dist. $\downarrow$ AUC $\uparrow$
Avg. Min. Avg. Min. Avg. Min.
  VAT None 0.230 0.161 0.835 0.202 0.133 0.869 0.182 0.116 0.892
VAT-GC Grad-CAM 0.210 0.144 0.848 0.188 0.121 0.884 0.173 0.107 0.898
Diffusion None 0.230 0.160 0.768 0.203 0.135 0.847 0.189 0.124 0.849
Diffusion-GC Grad-CAM 0.199 0.131 0.803 0.177 0.112 0.846 0.167 0.103 0.870
4.3.3 The Effect of the Grad-CAM Heatmaps
In this section, we test how much guidance signal is provided by the VQA priors. This cannot be directly tested in the semi-supervised scenario, so we tested by using the Grad-CAM heatmaps as a direct input to a fully supervised gaze following model. We train the teacher models of Sec. 4.2 using different amounts of labeled data only, and introduce a new baseline: Diffusion-GC, where the diffusion model trained on labeled images was used to refine the Grad-CAM heatmaps computed directly on the test set. Note that VAT-GC and Diffusion-GC require the Grad-CAM heatmaps of the test images during inference, which may not be feasible in an online setting as it necessitates access to a large vision-language model.
Tab. 4 presents the results. VAT-GC outperforms the baseline VAT model in all cases, and Diffusion-GC performs best in the distance metrics. Both the VAT and diffusion model show significant performance boosts when Grad-CAM heatmaps are directly input into the gaze following models, which offers evidence that the VQA Grad-CAM heatmaps bring strong prior knowledge. As illustrated in Sec. 4.2, the lower AUC for Diffusion and Diffusion-GC can be attributed to the AUC on GazeFollow being evaluated with group-level annotations and favoring heatmaps spanning larger areas. We analyzed this in more detail in the supplementary material.
4.4 Semi-supervised Gaze Following for Video
In this experiment, we finetune a pretrained image gaze following model to specific video scenes in a semi-supervised manner. We experiment on 10 TV shows from the VideoAttentionTarget dataset [9], comprising a total of 31,978 gaze annotations. For each video, we randomly select a clip that contains about 100 annotated frames (2–23% of the entire video, mostly below 10%) as groundtruth data. The remainder of the videos are regarded as unlabeled. Following [9], we started with the image version of the teacher model pre-trained on the GazeFollow dataset and extended it temporally through an additional ConvLSTM network. The model was then fine-tuned on the given annotations and the generated pseudo-annotations for each video. The experiment simulates a real-world scenario where only a few frames of a specific scene are annotated.
Fig. 6 presents the $L_{2}$-distance and AUC in video adaptation experiments on each of the 10 videos. Our method outperforms other pseudo-annotation generation methods in both metrics on almost all videos. When the baseline models do not perform well, the improvement of our method is more pronounced (2%–5% improvement in both metrics, more details in supplementary). These findings demonstrate that our method can be used effectively in video gaze following, needing to initially label only a few frames.
Figure 6: Applying gaze following to videos in a semi-supervised manner. GCDR achieves the best performance on almost all videos (Dist and AUC averaged across videos: GCDR: 0.142, 0.843; Semi-VAT-GC: 0.155, 0.818; Semi-VAT: 0.164, 0.797 ; VAT: 0.166, 0.773). More details in the Supplementary.
5 Conclusion
We have proposed the first approach for generating high-quality pseudo-labels for the semi-supervised gaze following task. We leverage the priors from large VL models by computing Grad-CAM heatmaps from a pretrained VQA model that is prompted with a gaze following question. The Grad-CAM heatmaps offer strong guidance to the gaze target, but can be noisy. This led to a novel diffusion-based refinement method that refines these initial pseudo annotations with an annotation prior. Our approach works well on both image and video gaze following tasks with significant savings in the annotation effort. We hope our method will lead to the collection of larger gaze following datasets with annotation effort similar to current datasets. We plan to apply diffusion-based refinement to “noisy” annotations generated by VL models, in new semi-supervised tasks.